OpenStudy (anonymous):

. Calculate the kinetic energy of a 100.0-kg meteor approaching the Earth at a speed of 10.0 km/s. Remember that 1 km = 1000 m.

OpenStudy (anonymous):

Well kinetic energy is given by the equation:\[K _{E}=\frac{ 1 }{ 2 }mv^2\] So plug in your mass and velocity for m and v: \[K _{E}=\frac{ 1 }{ 2 }(100kg)(10,000m/s)^2\] p.s. i converted 10 km to 10,000 meters Simplify:\[\frac{ 1 }{ 2 }(100kg)(100,000,000m^2/s^2)\] Further simplfy:\[\frac{ 1 }{ 2 }10,000,000,000\frac{ kg*m^2 }{s^2 }=\frac{ 1 }{ 2 }10,000,000,000J\] Final simplification:\[K _{E}=5,000,000,000 J=5,000,000kJ\] ANSWER : THE KINETIC ENERGY IS 5,000,000 kiloJoules

OpenStudy (anonymous):

WOW that meteor is fast see how the velocity affects the kinetic energy more than just mass because you square it to get the kinetic energy?